# Automatic Calibration of a Multi-Camera System with Limited Overlapping Fields of View for 3D Surgical Scene Reconstruction

### v2

@@ 1 Introduction
3D surgical scene reconstruction (3D-SSR) involves computing a 3D representation of all relevant entities during surgery, including the anatomy shape and texture, surgical instrument poses, and body and hand poses of the medical staff. It can facilitate error prevention, operative performance assessment, formative feedback by avoiding the need for manual review and assessment of surgical video [1, 2]. For tasks such as automatic workflow analysis and activity recognition, the use of geometric representations have recently been shown to outperform image-based representations while requiring less labeled data [3]. In surgical navigation and robotic surgery, 3D-SSR can help to reduce the sim-to-real gap by providing accurate and realistic 3D reconstructions of surgical scenes in which robots and machine learning (ML)-based applications can be trained before being deployed in the real world [4, 5, 6, 7]. Optical cameras remain the preferred solution for capturing 3D information related to medical staff, including their pose, surgeon hand movements, surgical instrument positions, and anatomy shape and texture. They are widely used for their ability to capture detailed visual information in a non-invasive manner while recording events in real-time. We focus on 3D-SRR based on a multi-camera system in this work.
@@
The first challenge is sparsely sampled viewpoints (also referred to as sparse-view or wide-baseline) which leads to failures of Structure-from-Motion (SfM) pipelines [8, 9] due to the lack of reliable point correspondences across views. The second challenge involves complex materials (glossy, metallic, and mirrored surfaces) and poorly textured surfaces in the OR which also defeat standard SfM pipelines. The use of easily identifiable markers with known geometry, placed on a manufactured planar calibration pattern (e.g., standard checkerboard, ChArUco [10], AprilTag [11], ArUco [12]), is often employed to address these technical limitations. Typically, a calibration pattern is moved by an operator within the volume intersection of the camera frustums, providing a large set of inter-view point correspondences. This family of methods is now established as the gold standard for calibrating multi-camera systems [13, 14]. Other methods, such as those utilizing manually moved LEDs or laser pointers [15, 16] as calibration objects, are also employed to obtain inter-view point correspondences. However, all these methods are time-consuming, operator-dependent and require a certain level of expertise. Such limitations make them impractical in busy environments like operating rooms, where time between surgeries is limited, and hospital staff very rarely, if ever, have the expertise required for computer vision or calibration techniques. The third challenge in calibrating a multi-camera system for 3D-SSR lies in the weak overlap between fields of view, particularly due to significant changes in camera location and zoom level relative to the size of the OR. For example, the small red region in Fig. 1, which represents the field of view of camera C3 (associated to the top right image) as seen by camera C1, corresponding to an appearing scale change of 17x, is a typical example of this issue. This problem arises from the diverse nature of objects to be reconstructed—ranging from wounds typically 5 to 30 cm in length to individuals moving within the OR, which can encompass up to 200 cubic meters—imposing high variations in camera positions and zoom levels. By limited overlapping fields of view, we refer in the sequel to limited overlaps arising from significant changes in image scale—caused by substantial variations in position, zoom level, or a combination of both. This differs from limited overlap due to diverging camera orientations, which is not inherently related to 3D-SSR and has been specifically addressed in the literature [17, 18]. In addition to being manual and operator-dependent, the aforementioned board-based calibration methods fail to address this third challenge. These methods rely on patterns with fixed feature sizes (e.g., checkerboards or planar markers), which makes it impossible to simultaneously detect features in cameras with significantly different locations and zoom levels (e.g., up to 12x in our experiments). One potential solution to this scale issue involves calibrating subsets of cameras, followed by a global registration. However, this approach necessitates multiple calibration boards with varying grid sizes and, in extreme cases, the addition of intermediate-scale cameras, thereby significantly increasing the complexity of the calibration process.
@@
Figure 1: Top: Examples of camera viewpoints used for 3D surgical scene reconstruction. From left to right: far-field camera C1 (ceiling-mounted GoPro), near-field camera C2 (surgical lamp-mounted GoPro), and near-field camera C3 (ceiling-mounted Canon CR-N300), zoomed in on the surgical field. Bottom: Inter-view point correspondences provided by our multi-camera calibration system, observed from the same viewpoints. From left to right: all correspondences seen by C1 (blue, green, and red), correspondences between C1 and C2 (green), and correspondences between C1 and C3 (red). The surgeon’s body and hand poses, represented by blue lines in the first row, are successfully reconstructed using the estimated camera poses.
@@
In contrast, we propose a novel method that achieves robust calibration without these additional requirements. Our method is fully automated, fast, accurate, and designed to address the challenges posed by 3D-SSR. The key idea of the method is the use of a ceiling-mounted projector that projects arrays of multi-scale markers (MSMs) onto the floor. An MSM consists of 2D patterns projected at varying scales, with centers that remain invariant to scale changes. This ensures that all projected markers are successfully extracted and accurately localized from all viewpoints, even when handling extreme variations in fields of view (see left and right in Fig. 1). Multiple arrays of MSMs are projected, providing a large number of uniformly distributed point correspondences across all views. The process is fully automated, operates with a standard entry-level projector, does not require projector calibration, and exploits static marker projections, ensuring highly accurate extraction of correspondences without facing motion blur or temporal synchronization issues. The method is fast, with a required projection duration of about 70 seconds, allowing for last-minute adjustments to camera positions, particularly those mounted on movable lamps. Our multi-camera calibration method accurately calibrates all cameras in our experiment with average reprojection errors of 0.28 pixel, while nicely dealing with large scale variations imposed by surgical settings. In contrast, off-the-shelf ChArUco [10], checkerboard-based calibration method [13, 14], and state-of-the-art Structure-from-Motion pipelines [8, 9] fail to recover all camera poses.
@@ 2 Methods
Figure 2: (a) Overview of the proposed setup. Our multi-camera calibration system employs an entry-level projector (P, projection in blue). Cameras are highlighted in red. In this example, GoPro cameras placed in far field and in near field on surgical lamps, and a zoomed in Canon CR-N300 (C3) are employed. (b) Images from a far-field camera (C4, top) and a near field camera (C3, bottom) during the projection of the proposed MSM. Top: MSM is successfully detected (green cross) in the far-field view for a large scale projection (left) and not detected otherwise (right). Bottom: MSM is not detected in the near-field view for a large scale projection (left) and successfully detected otherwise (green cross in the right image), therefore providing one point correspondence between C3 and C4 through their shared imaged center. This operation is repeated, providing a set of dense and uniformly distributed point correspondences.
@@ 2.1 Multi-scale markers
Euclidean pattern selection. We recall that our objective is to establish point correspondences between views. This places specific requirements on the choice of the Euclidean pattern $\mathcal{P}\in\left\{\phi:\mathbb{R}^{2}\to[0,1]\right\}$, which must meet two key criteria. The first criterion is that $\mathcal{P}$ allows for the computation of a point $\mathbf{c}$, referred to as its center, such that $\mathbf{c}=f(\mathcal{P})$, where $f$ denotes a detection function. The second criterion is that when $f$ is applied to an image of $\mathcal{P}$ transformed by any homography $g$, it returns the imaged center, namely $f(g(\mathcal{P}))=g(f(\mathcal{P}))=g(\mathbf{c})$. By choosing the marker center $\mathbf{p}$ such that $\mathbf{p}=\mathbf{c}$, this second criterion ensures the detection of any image of the marker center, since $f(g_{cam}(g_{proj}(h_{\mathbf{p}}(\mathcal{P})))=g(\mathbf{p})$, where $g_{\text{cam}}$ is the homography from the floor plane to the image plane of the camera. Among the patterns that satisfy these two criteria, a square is one example, where the imaged center can be computed using the intersection of its imaged diagonals, which is a projective invariant, or two (or more) concentric circles [19].
@@ 2.2 Calibration algorithm
We solve Eq. 4 in an incremental manner similarly to standard SfM pipelines [8]. The main difference lies in the initialization of the reconstruction, which consists of computing the relative camera motion of a well-chosen initial camera pair. Indeed, in our camera calibration scenario, correspondences are images of coplanar points which represent a degenerate case for estimating relative camera motion using standard epipolar geometry, specifically via the fundamental or essential matrix [20]. Instead, we compute the camera motion of the initial camera pair by decomposing the inter-image homography [20], which is the rigorous approach for handling coplanar or nearly coplanar points. The homography is estimated through RANSAC (using a threshold of 3 pixels) on normalized correspondences. The initial camera pair is selected based on the view score borrowed from the SfM pipeline [8]. This score represents the number and distribution of correspondences in a camera. We choose the initial camera pair as the one maximizing the score of both cameras among all possible camera pairs.
@@ 3.1 Calibration on synthetic data
We conducted a preliminary experiment to verify that the spatial distribution of 3D points located on the floor does not negatively impact calibration quality compared to the gold standard checkerboard calibration method [21, 13, 14], which uses 3D points within the volume intersection of the camera frustums, obtained using a moving calibration board.
@@
The first scenario, referred to as board volume, imitates the popular calibration approach with a hand-held ChArUco board placed randomly in front of the cameras. Simulated checkerboards are randomly distributed within a cylindrical working volume with a 3-meter radius and a maximum height of 1.5 meters. In the second scenario, referred to as board floor, the checkerboards are randomly distributed on the floor, within a circular working area on the floor of a 3-meter radius, simulating a calibration board moved on the floor. Both scenarios simulate a ChArUco board of size 120x85cm, which is identical to the one used in Section 3.2 and roughly corresponds to the minimal size that is detectable by all cameras in practical conditions. During the sampling of checkerboard poses, we ensure that the average number of observations per camera category is consistent across all scenarios, specifically around 2,000 for near-field and 3,000 for far-field cameras. The calibration is performed using the gold standard checkerboard method, which optimizes the board poses while constraining the points to align with the known board geometry.
@@
Evaluation The calibration quality is assessed by comparing the estimated camera poses with the ground truth. This is done by aligning the estimated point cloud to the ground truth using a Sim(3) similarity that minimizes the squared error.
@@
Figure 3: Rotation and translation RMSE over 200 simulations. See text for details.
@@ 3.2 Calibration on real-world data
Setup The experimental setup (see Fig. 2 (a)) included 9 cameras: 8 GoPro cameras (6 positioned in the far field and 2 mounted on surgical lamps near the surgical table in the near field) and 1 ceiling-mounted Canon camera with a strong optical zoom focusing on the surgical field. From this setup, we sampled different camera configurations to create four distinct setups with varying positions and viewpoints to test the robustness of the calibration method under setups of varying difficulty (Full, Easy, Medium, Hard). As we assume a (nearly) planar surface in our method, we removed the table during the calibration procedure. We compared our method to the gold standard multi-camera calibration method (same as used in Section 3.1) using a moving ChArUco board. We selected the ChArUco board with the smallest square size that was detectable in all cameras, including those in the far field. The board was manually moved throughout the scene to gather sufficient inter-view correspondences. In order to not penalize this method, the table was also removed during the calibration procedure. The calibration process took approximately 5 minutes to complete. Additionally, we evaluated the state-of-the-art SfM pipelines COLMAP [8] and GLOMAP [9] under two conditions: (i) in a fully furnished OR (including the table), and (ii) with a projected texture on the floor and the lights turned off to maximize contrast, with the OR table removed to avoid penalizing the method. For the proposed method, an entry-level projector (Acer H6518STi) was mounted on the OR ceiling. A sequence of 100 MSM arrays was projected onto the OR floor with the lights turned off. Each array contained 32 MSMs and was slightly offset from the previous one to create a dense grid of 3200 points across the floor. Seven scale factors, $\lambda\in\Lambda=\{1.0,1.4,2.0,3.0,4.0,6.0,8.0\}$, were used in the experiment, corresponding to marker sizes ranging from 5 cm to 40 cm. The experiment was repeated for MSMs generated from two different marker systems: AprilTags and ArUco. The proposed calibration procedure in our experiment took 70 seconds.
@@
Results All results presented for our method use ArUco markers and integrate the coplanarity constraint in the optimization. Results without the coplanarity constraint are very comparable and available in the supplementary material. Table 1 presents the reprojection errors for each distinct setup, while more detailed results for each camera category for the full setup are shown in Table 2. Table 3 evaluates the robustness by computing success rates for each camera setup, defined as the percentage of reconstructed cameras with a mean reprojection error below a threshold of 0.5/2/5 pixel on the evaluation dataset. We also compare the reconstructed point clouds and camera poses obtained using our method against those from the ChArUco baseline in the Full setup. The camera poses exhibit a rotational RMSE of 0.12 degrees and a translational RMSE of 6.13 mm, corresponding to a relative error of less than 0.14% of the mean inter-camera distance (4.22 m). The mean Euclidean error on the 3D points in the evaluation dataset is 1.63 $\pm$ 0.91 mm. Last, we show the accuracy of our calibration by applying it to 3D-SSR, as shown in Fig. 4. We successfully reconstruct the surgeon’s body and hands poses by locating the joints in the images and triangulating them using the cameras poses obtained with proposed calibration method. In that example, body joints were computed using [22], while hands joints were manually extracted.
@@
Table 1: Reprojection errors on calibration and evaluation data for the ChArUco-based and our proposed calibration method on distinct setups.
@@
Table 2: Results per camera category on the full setup for the ChArUco and our proposed method. Mean reprojection error (calib. & eval.), disribution score (calib., see Section 2.2), observations of 3D points with track length $>=2$ (calib.). The ChArUco-based method failed to register the close-up camera.
@@
Table 3: Success rates in % with a threshold of 0.5/2/5 pixel. See text for details.
Figure 4: Application of our calibration method to 3D-SSR. From left to right: reference image, the surgeon’s body poses, and hands, successfully reconstructed using camera poses computed with the proposed MSM-based calibration method. The reconstructed objects are integrated into the OR digital twin [23].

### v3

@@ 1 Introduction
3D surgical scene reconstruction (3D-SSR) involves computing a 3D representation of all relevant entities during surgery, including the anatomy shape and texture, surgical instrument poses, and body and hand poses of the medical staff. It can facilitate error prevention, operative performance assessment, formative feedback by avoiding the need for manual review and assessment of surgical video [MascagniASAMWAR22, RazaVDK22]. For tasks such as automatic workflow analysis and activity recognition, the use of geometric representations have recently been shown to outperform image-based representations while requiring less labeled data [pmlr-v227-hamoud24a]. In surgical navigation and robotic surgery, 3D-SSR can help to reduce the sim-to-real gap by providing accurate and realistic 3D reconstructions of surgical scenes in which robots and machine learning (ML)-based applications can be trained before being deployed in the real world [Gumbs21, Barnoyetal21, GumbsGBCSFIHPE22, JecklinJFFE22]. Optical cameras remain the preferred solution for capturing 3D information related to medical staff, including their pose, surgeon hand movements, surgical instrument positions, and anatomy shape and texture. They are widely used for their ability to capture detailed visual information in a non-invasive manner while recording events in real-time. We focus on 3D-SSR based on a multi-camera system in this work.
@@
The first challenge is sparsely sampled viewpoints (also referred to as sparse-view or wide-baseline) which leads to failures of Structure-from-Motion (SfM) pipelines [schonberger2016structure, pan2024glomap] due to the lack of reliable point correspondences across views. The second challenge involves complex materials (glossy, metallic, and mirrored surfaces) and poorly textured surfaces in the OR which also defeat standard SfM pipelines. The use of easily identifiable markers with known geometry, placed on a manufactured planar calibration pattern (e.g., standard checkerboard, ChArUco [garrido2014automatic], AprilTag [apriltags], ArUco [aruco]), is often employed to address these technical limitations. Typically, a calibration pattern is moved by an operator within the volume intersection of the camera frustums, providing a large set of inter-view point correspondences. This family of methods is now established as the gold standard for calibrating multi-camera systems [Bouguet2010, Rehder2016kalibr]. Other methods, such as those utilizing manually moved LEDs or laser pointers [led1, led2] as calibration objects, are also employed to obtain inter-view point correspondences. However, all these methods are time-consuming, operator-dependent and require a certain level of expertise. Such limitations make them impractical in busy environments like operating rooms, where time between surgeries is limited, and hospital staff very rarely, if ever, have the expertise required for computer vision or calibration techniques. The third challenge in calibrating a multi-camera system for 3D-SSR lies in the weak overlap between fields of view, particularly due to significant changes in camera location and zoom level relative to the size of the OR. For example, the small red region in Fig. 1, which represents the field of view of camera C3 (associated to the top right image) as seen by camera C1, corresponding to an appearing scale change of 17x, is a typical example of this issue. This problem arises from the diverse nature of objects to be reconstructed—ranging from wounds typically 5 to 30 cm in length to individuals moving within the OR, which can encompass up to 200 cubic meters—imposing high variations in camera positions and zoom levels. By limited overlapping fields of view, we refer in the sequel to limited overlaps arising from significant changes in image scale—caused by substantial variations in position, zoom level, or a combination of both. This differs from limited overlap due to diverging camera orientations, which is not inherently related to 3D-SSR and has been specifically addressed in the literature [simpleCalibNonOverlap, board3].
@@
Figure 1: Top: Examples of camera viewpoints used for 3D surgical scene reconstruction. From left to right: far-field camera C1 (ceiling-mounted GoPro), near-field camera C2 (surgical lamp-mounted GoPro), and near-field camera C3 (ceiling-mounted Canon CR-N300), zoomed in on the surgical field. Bottom: Inter-view point correspondences provided by our multi-camera calibration system, observed from the same viewpoints. From left to right: all correspondences seen by C1 (blue, green, and red), correspondences between C1 and C2 (green), and correspondences between C1 and C3 (red). The surgeon’s body and hand poses, represented by blue lines in the first row, are successfully reconstructed using the estimated camera poses
@@
We propose a novel method that is fully automated, fast, accurate, and designed to address the challenges posed by 3D-SSR. The key idea of the method is the use of a ceiling-mounted projector that projects arrays of multi-scale markers (MSMs) onto the floor. An MSM consists of 2D patterns projected at varying scales, with centers that remain invariant to scale changes. This ensures that all projected markers are successfully extracted and accurately localized from all viewpoints, even when handling extreme variations in fields of view (see left and right in Fig. 1). Multiple arrays of MSMs are projected, providing a large number of uniformly distributed point correspondences across all views. The process is fully automated, operates with a standard entry-level projector, does not require projector calibration, and exploits static marker projections, ensuring highly accurate extraction of correspondences without facing motion blur or temporal synchronization issues. The method is fast, with a required projection duration of about 70 seconds, allowing for last-minute adjustments to camera positions, particularly those mounted on movable lamps. Our multi-camera calibration method accurately calibrates all cameras in our experiment with average reprojection errors of 0.28 pixel, while nicely dealing with large scale variations imposed by surgical settings. In contrast, off-the-shelf ChArUco [garrido2014automatic], checkerboard-based calibration method [Bouguet2010, Rehder2016kalibr], and state-of-the-art SfM pipelines [schonberger2016structure, pan2024glomap] fail to recover all camera poses. The source code is publicly available at https://github.com/tflueckiger/calib-proj.
Figure 2: (a) Overview of the proposed setup. Our multi-camera calibration system employs an entry-level projector (P, projection in blue). Cameras are highlighted in red. In this example, GoPro cameras placed in far field and in near field on surgical lamps, and a zoomed in Canon CR-N300 (C3) are employed. (b) Images from a far-field camera (C4, top) and a near field camera (C3, bottom) during the projection of the proposed MSM. Top: MSM is successfully detected (green cross) in the far-field view for a large scale projection (left) and not detected otherwise (right). Bottom: MSM is not detected in the near-field view for a large scale projection (left) and successfully detected otherwise (green cross in the right image), therefore providing one point correspondence between C3 and C4 through their shared imaged center. This operation is repeated, providing a set of dense and uniformly distributed point correspondences
@@ 2.1 Multi-scale markers
Euclidean pattern selection We recall that our objective is to establish point correspondences between views. This places specific requirements on the choice of the Euclidean pattern $\mathcal{P}\in\left\{\phi:\mathbb{R}^{2}\to[0,1]\right\}$, which must meet two key criteria. The first criterion is that $\mathcal{P}$ allows for the computation of a point $\mathbf{c}$, referred to as its center, such that $\mathbf{c}=f(\mathcal{P})$, where $f$ denotes a detection function. The second criterion is that when $f$ is applied to an image of $\mathcal{P}$ transformed by any homography $g$, it returns the imaged center, namely $f(g(\mathcal{P}))=g(f(\mathcal{P}))=g(\mathbf{c})$. By choosing the marker center $\mathbf{p}$ such that $\mathbf{p}=\mathbf{c}$, this second criterion ensures the detection of any image of the marker center, since $f(g_{cam}(g_{proj}(h_{\mathbf{p}}(\mathcal{P})))=g(\mathbf{p})$, where $g_{\text{cam}}$ is the homography from the floor plane to the image plane of the camera. Among the patterns that satisfy these two criteria, a square is one example, where the imaged center can be computed using the intersection of its imaged diagonals, which is a projective invariant, or two (or more) concentric circles [concentriccircles].
@@ 2.2 Calibration algorithm
We solve Eq. 4 in an incremental manner similarly to standard SfM pipelines [schonberger2016structure]. The main difference lies in the initialization of the reconstruction, which consists of computing the relative camera motion of a well-chosen initial camera pair.
Indeed, in our camera calibration scenario, correspondences are images of coplanar points which represent a degenerate case for estimating relative camera motion using standard epipolar geometry, specifically via the fundamental or essential matrix [Hartley2003MVG]. Instead, we compute the camera motion of the initial camera pair by decomposing the inter-image homography [Hartley2003MVG], which is the rigorous approach for handling coplanar or nearly coplanar points. The homography is estimated through RANSAC (using a threshold of 3 pixels) on normalized correspondences.
The initial camera pair is selected based on the view score borrowed from the SfM pipeline [schonberger2016structure]. This score represents the number and distribution of correspondences in a camera. We choose the initial camera pair as the one maximizing the score of both cameras among all possible camera pairs.
@@ 3.1 Calibration on synthetic data
We conducted a preliminary experiment to verify that the spatial distribution of 3D points located on the floor does not negatively impact calibration quality compared to the gold standard checkerboard calibration method [Zhang2000CameraCalibration, Bouguet2010, Rehder2016kalibr], which uses 3D points within the volume intersection of the camera frustums, obtained using a moving calibration board.
@@
The first scenario, referred to as board volume, imitates the popular calibration approach with a hand-held ChArUco board placed randomly in front of the cameras. Simulated checkerboards are randomly distributed within a cylindrical working volume with a 3-meter radius and a maximum height of 1.5 meters.
In the second scenario, referred to as board floor, the checkerboards are randomly distributed on the floor, within a circular working area on the floor of a 3-meter radius, simulating a calibration board moved on the floor. Both scenarios simulate a ChArUco board of size 120x85cm, which is identical to the one used in Section 3.2 and roughly corresponds to the minimal size that is detectable by all cameras in practical conditions. During the sampling of checkerboard poses, we ensure that the average number of observations per camera category is consistent across all scenarios, specifically around 2,000 for near-field and 3,000 for far-field cameras. The calibration is performed using the gold standard checkerboard method, which optimizes the board poses while constraining the points to align with the known board geometry.
@@
Evaluation The calibration quality is assessed by comparing the estimated camera poses with the ground truth. This is done by aligning the estimated point cloud to the ground truth using a Sim(3) similarity that minimizes the squared error via Umeyama’s algorithm [umeyama].
@@
Figure 3: (a) Rotation RMSE and (b) translation RMSE over 200 simulations. See text for details
@@ 3.2 Calibration on real-world data
Setup The experimental setup (see Fig. 2 (a)) included 9 cameras: 8 GoPro cameras (6 positioned in the far field and 2 mounted on surgical lamps near the surgical table in the near field) and 1 ceiling-mounted Canon camera with a strong optical zoom focusing on the surgical field. From this setup, we sampled different camera configurations to create four distinct setups with varying positions and viewpoints to test the robustness of the calibration method under setups of varying difficulty (Full, Easy, Medium, Hard). As we assume a (nearly) planar surface in our method, we removed the table during the calibration procedure. We compared our method to the gold standard multi-camera calibration method (same as used in Section 3.1) using a moving ChArUco board. We selected the ChArUco board with the smallest square size that was detectable in all cameras, including those in the far field. The board was manually moved throughout the scene to gather sufficient inter-view correspondences. In order to not penalize this method, the table was also removed during the calibration procedure. The calibration process took approximately 5 minutes to complete. Additionally, we evaluated the state-of-the-art SfM pipelines COLMAP [schonberger2016structure] and GLOMAP [pan2024glomap] under two conditions: (i) in a fully furnished OR (including the table), and (ii) with a projected texture on the floor and the lights turned off to maximize contrast, with the OR table removed to avoid penalizing the method. For the proposed method, an entry-level projector (Acer H6518STi) was mounted on the OR ceiling. A sequence of 100 MSM arrays was projected onto the OR floor with the lights turned off. Each array contained 32 MSMs and was slightly offset from the previous one to create a dense grid of 3200 points across the floor. Seven scale factors, $\lambda\in\Lambda=\{1.0,1.4,2.0,3.0,4.0,6.0,8.0\}$, were used in the experiment, corresponding to marker sizes ranging from 5 cm to 40 cm. The experiment was repeated for MSMs generated from two different marker systems: AprilTags and ArUco. The proposed calibration procedure in our experiment took 70 seconds.
@@
Results All results presented for our method use ArUco markers and integrate the coplanarity constraint in the optimization. Results without the coplanarity constraint are very comparable and available in Online Resource 1. Table 1 presents the reprojection errors for each distinct setup, while more detailed results for each camera category for the full setup are shown in Table 2. Table 3 evaluates the robustness by computing success rates for each camera setup, defined as the percentage of reconstructed cameras with a mean reprojection error below a threshold of 0.5/2/5 pixel on the evaluation dataset. We also compare the reconstructed point clouds and camera poses obtained using our method against those from the ChArUco baseline in the Full setup. The camera poses exhibit a rotational RMSE of 0.12 degrees and a translational RMSE of 6.13 mm, corresponding to a relative error of less than 0.14% of the mean inter-camera distance (4.22 m). The mean Euclidean error on the 3D points in the evaluation dataset is 1.63 $\pm$ 0.91 mm. Last, we show the accuracy of our calibration by applying it to 3D-SSR, as shown in Fig. 4. We successfully reconstruct the surgeon’s body and hands poses by locating the joints in the images and triangulating them using the cameras poses obtained with proposed calibration method. In that example, body joints were computed using [openpose], while hands joints were manually extracted.
@@
Table 1: Reprojection errors on calibration and evaluation data for the ChArUco-based and our proposed calibration method on distinct setups
@@
Table 2: Results per camera category on the full setup for the ChArUco and our proposed method. Mean reprojection error (calib. & eval.), disribution score (calibration data, see Section 2.2), observations of 3D points with track length >=2>=2 (calibration data). The ChArUco-based method failed to register the close-up camera
@@
Table 3: Success rates in % with a threshold of 0.5/2/5 pixel. See text for details
Figure 4: Application of our calibration method to 3D-SSR. From left to right: reference image, the surgeon’s body poses, and hands, successfully reconstructed using camera poses computed with the proposed MSM-based calibration method. The reconstructed objects are integrated into the OR digital twin [Hein_2024_CVPR]. An animation is available in Online Resource 2